APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday August 27, 2001 (MJD 52149) ---=== OBSERVING TEAM ===--- swing: Steph Snedden night: Mike Harvanek Jurek Krzesinski ---=== OBSERVING PLAN ===--- Spectroscopy, and after moon set imaging. ---=== OBSERVING SUMMARY ===--- One plate finished (666). Partial data for 2 plates (634, 644). 18:50 telescope checked/ready, doors open, the sky mostly clear. 19:50 enclosure off cart 2 plt 234 on the telescope. Thin clouds. 20:50 On sky - we took 4 science exposure at miserable seeing -got very low S/N (below 2). 22:45 Cart 9 plt 644 on. 5 science exposure taken but still with a bad S/N ratio. 00:40 telescope closed - humidity too high. 04:00 enclosure off, catr 3 plt 666 mounted, we are back on the sky. 05:10 We finished that plate and declared it done. 05:50 telescope in the enclosure: cart 2 plt 634 on. 05:50 dewars forced to fill up, hoses unhooked!, wind baffle off and locked out, telescope tracking stow position ************************************* WE ARE LEAVING TELESCOPE TRACKING !!! track 121,30 mount /rota=0 /rotty=mount WIND BAFFLE off and locked, ************************************* Problems: the same as last night: 1) one or more axes did not end on a full step 0, 0, 0, 0, 1" 2) dsc_I_trcInfo Applying correction 3859286 to azimuth ---=== OBSERVING LOG ===--- 18:50 Ready to open. Some clouds at the Sun set but moustly blue sky. Cart 2 plt 634 on the telescope. 20:05 Ready to go. 20:50 An fk5 star found with no problem using the last night offsets. Again the seeing is very bad from 2.5 - 4.0", 3.5 m telescope observrs reported seeing about 4". The efficiency read from the guider: ~25%. There are thin clouds around, and we are collecting noise, so we decided do postcalib and no smearing. We took 4 science exposures. The sky was too bright for all of them. Only one of these exposures had (S/N)^2 > 2 and this was for only one camera (r2). Son-of-spectro complained about red monsters, whopping fibers, scattered light and large flexure. Again we have problems with "fiducial reading" first time occured: Aug 27 22:34:37 sdssmcp tLatch dsc_I_trcInfo Applying correction 3859286 to azimuth Aug 27 22:34:37 sdssmcp tLatch dsc_E_trcErr correction for azimuth 3859286 is too large Aug 27 22:34:37 sdssmcp tLatch dsc_W_trcWarn Not disabling MS.ON; error 3859286 (max allowed: 600) and later. And we keep ignoring: 0 0 I Received="?GOCORR one or more axes did not end on a full step 0, 0, 0, 0, 1" 22:45 Cart 9 plt 644 on. We didn't find stars with go to field, but we've found quickly fk5 star in a fiber. The seeing is getting better! However wind and its speed remain the same... The efficiency is high over 60%, but we've got very low S/N. Perhaps the sky brightness is too high? We did 5 science exposures for this plate. The sky was too bright for all of them. Three of the exposures had (S/N)^2 > 2 for r2. No useful data for b2 and r1. Numbers were not available for b1 because son-of-spectro couldn't reduce the arc but it is doubtful any of the data were usable. Son-of spectro again complained about whopping fibers, scattered light and large flexure. 23:30 Thick clouds are coming. 00:30 We finished exposure and postcalib (needed because we SOP couldn't reduce data in blue 1). 00:40 High humidity- closing. 03:30 Humidity rapidly dropped down to 60% , we waited a while and reopened the telescope, with cart 3 plt 666 on. 04:00 enclosure off, catr 3 plt 666 mounted, we are back on the sky. Seeing 2.5", no moon and the efficiency over 60%. Bizarre - Mike in his SAOimage window couldn't see a star in the guider 11-th fiber, but Jurek in his control SAOimage window saw stars in all fibers - can it be explained (cart 3 plt 666)? 04:40 We've finished cart 3 plt 666 with a first exposure (keep in mind that last night we've got S/N > 12 for that plate). Since there is no more plates available at this time, we decided to take a few more exposures for this plate. 05:20 Mounting cart into 2 plt 634 and closing the telescope. 05:50 telescope in enclosure: cart 2 plt 634 on. 05:50 dewars forced to fill up, hoses unhooked! WE ARE LEAVING TELESCOPE TRACKING !!! track 121,30 mount /rota=0 /rotty=mount WIND BAFFLE off and locked, endNight -doSpectro -doTpm : data written without any problems. Landru: all windows open, going... ---=== IMAGING RUN SUMMARY ===--- 2516 100 O ignore 68.77 77.39 68 ---=== IMAGING RUN DETAILS ===--- Run Stripe Strip Flavor Lambda Last Frame # Begin End ------------------------------------------------------------------ 2515 100 O bias -4.57 30.13 30 test (doghouse) 2516 100 O ignore 68.77 77.39 68 test (doghouse) ---=== SKIPPY RESULTS ===--- ---=== LTMATCH RESULTS ===--- ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment =============================================== ----- sequence 10766, plate -9999 ------- 20:41 10766 0.0 bias {{}} 23:48 10767 0.0 bias {{}} ----- sequence 10768, plate 634 ------- 03:07 10768 10.0 flat pre-flat 03:09 10769 2.0 arc pre-arc 03:33 10770 900.0 target science 03:51 10771 901.0 target science 04:08 10772 901.0 target science 04:28 10773 900.0 target science 04:31 10774 10.0 flat post-flat 04:33 10775 2.0 arc post-arc Plate 634 was observed in sequence(s) 10768 for a total of 3602 seconds ----- sequence 10776, plate 644 ------- 04:55 10776 10.0 flat pre-flat 04:58 10777 2.0 arc pre-arc 05:18 10778 901.0 target science 05:35 10779 900.0 target science 05:53 10780 900.0 target science 06:11 10781 900.0 target science 06:31 10782 900.0 target science 06:36 10783 10.0 flat post-flat 06:38 10784 2.0 arc post-arc Plate 644 was observed in sequence(s) 10776 for a total of 4501 seconds ----- sequence 10785, plate 666 ------- 10:05 10785 10.0 flat pre-flat 10:07 10786 2.0 arc pre-arc 10:26 10787 901.0 target science 10:44 10788 900.0 target science 11:02 10789 900.0 target science 11:09 10790 240.0 smear {{Smear Pattern}} 11:12 10791 10.0 flat post-flat 11:15 10792 2.0 arc post-arc Plate 666 was observed in sequence(s) 10785 for a total of 2701 seconds ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Cart Plate getclock Az Alt Rot Scale 21:18 2 634 998968690 0.000068 0.001982 0.011646 1.000000 23:03 9 644 998974988 -0.000899 -0.000016 0.018786 1.000000 04:13 3 666 998993597 -0.000998 -0.000265 0.027669 0.999950 04:23 3 666 998994212 -0.000998 -0.000265 0.027669 0.999890 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL4344 Stays: JL4345 ---=== FOCUS LOG ===--- Time dpTemp dewPt Press TmpOut HumOut Windd Winds 21:09 55.6 41.4 21.60 58 71.8 19.0 13 setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm 21:10 2 634 1.00000 0 1113 -100 44 40.8 13.1 19 72 gdr 4.0 23:02 9 644 1.00000 0 1100 -100 30 44.5 12.2 22 73 gdr 2.0 00:09 9 644 1.00000 0 1084 -100 7 49.0 11.9 18 72 gdr 4.0 04:10 3 666 0.99995 493 1497 -100 10 47.7 11.9 3 9 gdr 2.5 04:24 3 666 0.99989 1085 1836 -250 5 48.1 12.1 10 81 gdr 2.1 ---=== WEATHER LOG ===--- 19:52 58 41 19 73 (ENE) 178 - 20:25 56 41 23 71 (ENE) 217 - 20:57 56 41 24 75 (ENE) 265 - 21:13 55 41 15 78 (ENE) 251 - 21:29 55 41 19 76 (ENE) 264 - 22:00 55 41 19 73 (ENE) 242 - 22:31 54 42 21 76 (ENE) 231 - 23:02 53 43 20 82 (E) 234 - 23:32 53 44 15 17 (NNE) 239 - 00:02 54 45 19 74 (ENE) 249 - 00:32 52 47 9 86 (E) 238 - 01:55 51 48 17 70 (ENE) 217 - 02:25 52 47 13 65 (ENE) 243 - 02:55 51 47 15 54 (NE) 184 - 03:27 54 37 3 351 (NNW) 186 - 03:58 53 38 4 359 (NNW) 177 - 04:06 53 40 4 47 (NE) 235 - 04:36 52 40 12 69 (ENE) 213 - 05:09 52 42 12 69 (ENE) 194 - 05:39 52 40 7 65 (ENE) 167 - 06:11 52 41 6 62 (ENE) 194 - ---=== TELESCOPE STATUS ===--- 18:50 telescope checked/ready, doors open, cart 2 634 mounted. 19:50 enclosure off. 20:05 ready to observe. 00:40 telescope closed - humidity too high. 04:00 enclosure off, catr 3 plt 666 mounted, we are back on the sky. 05:50 telescope in enclosure: cart 2 plt 634 on, dewars forced to fill, WIND BAFFLE off and locked, tracking 121,30 mount /rota=0 /rotty=mount ---=== SOFTWARE USED ===--- IOP/SOP: v3_53_7 Watcher: v2_9_0 MCP: v5_3_1 TPM: v2_4_1 AstroDa: v14_36 TCC: 2.5.5 sdssProcedures: v1_22 tccMon : v1_6 ---=== MIRROR NUMBERS ===--- Time dpTemp dewPt Press TmpOut HumOut Windd Winds 19:18 59.9 41.1 21.59 49 75.0 18.1 15 PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: 0.0710 0.0710 Axial B: -0.0890 -0.0980 Axial C: 0.0640 0.0610 Trans D: 4.1850 4.1800 Lateral E: 1.9340 1.9365 Lateral F: 0.4345 0.4370 GALILS Commanded: 5377. -3699. 923. -208. 31535. 30647. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES primDesOrient: 0.00 -11.80 23.00 1256.90 642.10 primOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.0 Air temp: 15.5 C Altitude: 30.09 MIGS TONIGHT NOMINAL Axial A: -1.2950 -1.2900 Axial B: -1.3040 -1.3150 Axial C: -1.3140 -1.3060 Trans D: 0.1360 0.1890 GALILS Commanded: 1583850. 1583850. 1583850. 2600. 3800. Actual: 1583851. 1583857. 1583825. 2600. 3800. SETMIR VALUES secDesOrient: 1257.00 0.00 0.00 0.00 -175.00 secOrient: 1257.01 0.01 0.00 0.00 -175.13 ---=== PROBLEMS IN DETAIL ===---